OpenStudy (anonymous):

Write an expression for the phrase. The product of x and negative 5.

OpenStudy (solomonzelman):

when you say product of x and 3, you would write \(\normalsize\color{black}{ \rm {x\times 3 }}\). when you say product of 4 and n, you would write \(\normalsize\color{black}{ \rm {4\times n }}\). when you say product of (-4) and v, you would write \(\normalsize\color{black}{ \rm {(-4)\times v }}\). Can you do your's now ?

OpenStudy (anonymous):

x * (-5) i have the same math as you

OpenStudy (solomonzelman):

Yes ─────────────────────────── ─────────────────────────── 1) Click and hold ALT 2) click the number code (using the numbers that are on the right of the keyboard, and `NOT` the ones below `F1`, `F2`, `F3`, etc., ) 3) release the ALT number code result ` 2 5 1 ` √ `0 2 1 5 ` × `2 4 6 ` ÷ `7 5 4 ` ≥ `7 5 5` ≤ `2 4 1 ` or `7 5 3` ± `2 4 7` ≈ `0 1 8 5 ` ¹ `2 5 3 ` ² `0 1 7 9 ` ³ if you want to be able to put these symbols without copy pasting at all times.

OpenStudy (anonymous):

is it the the same as (-5) * x

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

But in your case you just need to write the expression, and it would sound the same way, despite that IT IS INDEED the same thing. Just write x × (-5) .

OpenStudy (solomonzelman):

it *wouldn't sound

OpenStudy (anonymous):

thank you